OpenStudy (anonymous):

find the solution for 2x^2+11x+5=0

Parth (parthkohli):

First, Factor the terms. \(\Large \color{midnightblue}{\rightarrow 2x^2 + 10x + x + 5 = 0 }\) \(\Large \color{midnightblue}{\rightarrow 2x(x + 5) + 1(x + 5) = 0 }\) \(\Large \color{midnightblue}{\rightarrow (2x + 1)(x + 5) = 0 }\) Either 2x + 1 = 0 or x + 5 = 0 2x = -1 x = -1/2 OR x + 5 = 0 x = -5

OpenStudy (ujjwal):

http://www.purplemath.com/modules/quadform.htm